# Data-Juicer 2.0: Cloud-Scale Adaptive Data Processing for Foundation Models

### v1

@@ 4.3. Web & CMD Tools
Thanks to the low-level and RESTful APIs, we further establish service-level capabilities across various scenarios in Data-Juicer 2.0, encapsulating scenario-specific processing solutions into built-in data recipes, described as YAML configuration files for end-to-end workflows (an example for video-data-synthesis is shown in Fig. 4). Centered around these data recipes, we provide user-friendly, highly encapsulated command-line interfaces as illustrated in Listing 4.3, supporting automated lightweight installation and recipe-level invocation. These tools automatically gather and install the requisite of OPs within the recipe, perform automated OP fusion, and maximize adaptive utilization of computing resources while minimizing user cognitive load and operational cost.
@@
Utilizing the aforementioned APIs and tools, Data-Juicer’s capabilities are seamlessly integrated into Alibaba Cloud’s visual modeling product, PAI-Designer (Cloud, 2024d). As shown in Fig. 1, DJ-Operators are encapsulated as visual components within Designer, presented in a directory tree structure for easy access. Users can manage data by intuitive drag-and-drop manners, and import built-in or customized recipes in YAML format using graphical canvas interface. With Designer as entry point, users can activate MaxCompute cluster nodes (Cloud, 2024c) or containerized deep compute clusters (PAI-DLC) (Cloud, 2024b) directly for efficient cloud-native distributed data processing, leveraging Alibaba Cloud’s data storage solutions (e.g., Object Storage OSS (Cloud, 2024f), Cloud Parallel File System CPFS (Cloud, 2024a), and NAS (Cloud, 2024e)) along with automated performance monitoring and optimization.
4.4. Natural Language Interaction
User:: Filter out too short text samples. The dataset path is input_data.jsonl Assistant:: <thought>Filter out text samples that are too short based on the provided dataset path.</thought>
@@ 4.4. Natural Language Interaction
Assistant:: <thought>The filtering has been successfully executed, and the filtered dataset is now at the specified path.</thought>
@@
Assistant:: Your request to filter out too short text samples has been completed. The processed dataset is available at output.jsonl.
Querying and understanding over 100 diverse OPs have become challenging. Thankfully, advances in foundation models have transformed human-computer interactions. Research shows the synergy of reasoning and acting (ReAct) (Yao et al., 2022) in language models enhances task-solving. Using RESTful API services, we developed model-driven agents based on ReAct, enabling intuitive data processing through natural language. We adopted AgentScope (Gao et al., 2024), a multi-agent platform, for low-code integration, using prompt-optimized ReAct agents to align OP functionalities with our RESTful APIs. Built-in function preprocessing, response parsing, and fault-tolerance streamline development. Data-Juicer 2.0 thus allows flexible instruction input, interpreting user intents to execute appropriate data processing functions.
@@
5. Data-Juicer Core Runtime
5.1. Data-Juicer-Dataset
5.1.1. Programming Interfaces with Extensive Compatibility
The core runtime of Data-Juicer 2.0 is implemented in Python, featuring a flexible data processing pipeline characterized by two execution modes: a standalone mode for convenient single-node execution, based on the Hugging Face Dataset (Lhoest et al., 2021), and a distributed mode that offers scalability across multiple nodes, leveraging the Ray Data (Moritz et al., 2018) and MaxFrame-DataFrame (Cloud, 2024c). These frameworks offer diverse computational engines, each with distinct capabilities and suitable use cases, but they also come with heterogeneous programming interfaces and complex implementation details.
In Data-Juicer 2.0, we exploit the strengths of these diverse computational engines and their associated programming classes by abstracting a top-level DJ-Dataset class and devising a standardized data representation schema (detailed in the next subsection). The primary design principle employed is the Facade Pattern. The abstracted class provides comprehensive and unified development interfaces, facilitating seamless use across standalone, distributed, and cloud cluster environments, meanwhile optimized to support extensive multimodal data processing.
@@ 5.1.1. Programming Interfaces with Extensive Compatibility
data.export(tgt_file) # tgt_file can be either local path or remote path Some basic interfaces of DJ-Dataset.
Moreover, following the principle of Template Method Pattern, the DJ-Dataset class establishes templated workflows for OP executing on datasets and provides several basic and unified interfaces, catering to users with diverse programming preferences, as demonstrated in Listing 5.1.1. The complexities of underlying computational engines and cumbersome runtime issues related to instantiated dataset objects are abstracted away from these interfaces, ensuring transparency and ease of use for end-users and developers alike.
New functionalities can be added by merely implementing new OP classes without altering the internal logic of DJ-Dataset. The standalone and distributed execution modes automatically recognize and switch configurations based on runtime settings. Coupled with unified function signatures, the data processing can be consolidated into a batched processing mode, not only standardizing the data interface for OP processing but also facilitating the implementation of a robust sample-level fault tolerance mechanism. These internal adaptation features will be elaborated in Sec. 5.4.2.
@@
[!h] {minted} [ frame=lines, framesep=2mm, baselinestretch=1.2, fontsize=, linenos, breaklines=true, bgcolor=lightgray!10, rulecolor=black, numbersep=5pt, ] json ”text”: ”¡__dj__image¿Antarctica is Earth’s southernmost & least-populated ” ”continent. ¡—__dj__eoc—¿¡__dj__video¿¡__dj__audio¿ Situated almost ” ”entirely south of the Antarctic Circle, it contains the geographic ” ”South Pole. ¡—__dj__eoc—¿Antarctica is the 5th-largest continent, ” ”about 40”Most of Antarctica is covered by the ice sheet. ¡—__dj__eoc—¿”, ”images”: [ ”path/to/the/image/of/antarctica_snowfield”, ”path/to/the/image/of/europe_map” ], ”audios”: [”path/to/the/audio/of/sound_of_waves_in_Antarctic_Ocean”], ”videos”: [”path/to/the/video/of/remote_sensing_view_of_antarctica”], ”meta”: ”version”: ”0.1”, ”text_category_tags”: ”xxx” , ”stats”: ”lang”: ”en”, ”image_heights”: [336, 728], ”audio_duration”: [5], ”video_motion_score”: [1.3], Illustration of data schema of Data-Juicer 2.0.
5.1.2. Flexible Data Schema
@@
In the context of foundation models, datasets are complex, originating from various sources and modalities. Target tasks are diverse, requiring different data formats and organizational structures. Integrating diverse data processing into a unified pipeline is challenging. Data-Juicer 2.0 adopts an extensible data representation schema that intrinsically supports multimodal data types and flexible data field customization.
@@
An example of data sample is illustrated in Listing 5.1.1. In the data schema, each DJ-Dataset object encapsulates column-stored samples represented by a non-recursive dictionary whose level-1 fields include three categories: (1) the data payload within customizable fields such as text, images, and audios; (2) the meta information concerning this sample that stems from either its raw data or Data-Juicer’s Mappers for tagging; (3) the stats information calculated by Data-Juicer’s Filters (all Filter OPs consist of a compute_stats function, utilized for subsequent analysis and conditional filtering).
@@
For the data payload fields, the content is primarily centered on the textual modality within the text field, capable of optionally including several text chunks split by customizable end-of-chunk tokens, which are <|__dj__eoc|> in default. Each chunk serves as a semantic unit, where the associated multimodal data within the same chunk relates to the same topic, thereby aligning with one another. Multimodal data (excluding text) are denoted by ordered customized special tokens in the text (e.g., <__dj__video>, <__dj__image>), storing entities at designated files, accessible through indexed file paths within the modality fields of samples to facilitate object sharing and lazy evaluate.
For instance, in the example shown in Listing 5.1.1, there are four chunks split by four <__dj__eoc> tokens in the text field, and two images represented by <__dj__image> tokens and stored in the images field. The former image is located in the 1st chunk and the latter image is in the 3rd chunk. All of them are visual supplementary to the textual contents in the corresponding chunks.
@@
This token-centric representation elegantly aligns multimodal data while preserving positional information, friendly to current prevalent learning paradigms for foundation models employing next-token prediction tasks (Bai et al., 2024b; Wang et al., 2024), and beneficial for data storage—locally, data samples are represented as-is, globally, identical portions across distinct data samples are redundantly stored once.
@@
Also note the Data-Juicer data schema’s compatibility with simple cross-modal pairing datasets (e.g., image-text pairs (Chen et al., 2015) and video-text pairs (Xu et al., 2023)), as well as complex interleaved multimodal datasets (e.g., MMC4 (Zhu et al., 2023)), owing to its chunkable and special token design. Additionally, we undertook substantial efforts to ensure Data-Juicer’s compatibility with the foundation model community, such as aligning post-tuning related fields (e.g., query, response) with the ModelScope-SWIFT repository (Zhao et al., 2024) for smooth training of 400+ foundation models. A suite of dataset conversion tools was also developed to efficiently transform various popular multimodal datasets to and from the Data-Juicer schema, enabling users to process extensive existing models and datasets within Data-Juicer. These tools also serve as demonstrations for extending support to other uncovered datasets.
5.2. Data-Juicer-Operators
5.2.1. Operator Types and Design Principles.
@@ 5.2. Data-Juicer-Operators
Operators (OPs) are fundamental units responsible for executing processing functionalities, such as enhancing the accuracy of descriptive text or filtering out images containing Not-Safe-for-Work (NSFW) content. As illustrated by the middle orange box in Fig. 1, Data-Juicer defines five previous atomic OP classes: Formatter, Filter, Mapper, Deduplicator, and Selector; alongside four new types of compositional OPs: Grouper, Aggregator, FusedOP, and ScriptOP. The first five OP classes handle dataset format conversion, sample filtering, modification, deduplication, and selection, respectively. Following the Strategy Pattern, they are designed to encapsulate diverse algorithms that can be dynamically used to process the data. Each OP has a clearly defined role and can be interchanged or modified without impacting other system parts.
@@
Moreover, following the Decorator Pattern, compositional OPs are provided to enhance existing functionality without modifying the prevailing OP structure while dynamically adding data processing behavior to objects. The FusedOP enables explicitly grouping multiple atomic OPs to process data in a fine-grained manner within the same data batch, opposed to the default sequential processing across datasets as demonstrated by Listing 5.2.1. The Grouper takes a DJ-Dataset as input and groups data samples into batches, which can then be input into the Aggregator for subsequent aggregation. For example, we can employ extract_entity_attribute_mapper and entity_attribute_aggregator following a key_value_grouper for meta-information extraction from textual input. Meanwhile, the ScriptOP allows users to incorporate existing Python files or execute code snippets, utilizing customized functionalities encoded within scripts like helper_func.py. This includes leveraging existing Data-Juicer command tools or integrating short Python scripts (e.g., lambda functions).
@@
5.2.2. Unified Coordination of Logical Operations.
@@
Specifically, a top-level OP factory class is abstracted above the aforementioned fundamental OP classes. In this class, functionalities common to all OPs are extracted, such as preprocessing of instantiated parameters, support for serialization, and configuration of OP-aware runtime parallelism. Besides, a unified run() method is implemented, maintaining a consistent interface for integration and API calls. Furthermore, parallelism in multi-processing and multi-GPU is automatically configured and decoupled from specific OPs, ensuring transparency for end users and developers, as introduced subsequently in Sec. 5.5. Lower-level OP classes define their own templated execution logic behind the run invocation. Taking the Filter class as an example, its core logic first engages the compute_stats() method to obtain statistical information based on specific metrics and then invokes its process() method to determine sample filtering based on thresholds.
@@ 5.3. Processing Job Control
With the fundamental DJ-Dataset and DJ-Operators primitives established, we offer a series of control panel modules (the red box in Fig. 1) to organically combine them and accomplish data processing tasks. Executor encapsulates a series of standardized execution processes tailored for different standalone and distributed engines, leveraging modules such as Config and Monitor to accomplish end-to-end system configuration, data loading, analysis, processing iteration, data checkpointing, and more. Template workflows manage the complete data development lifecycle, including feedback, data storage, logging, and performance and operational monitoring.
The processing capability of Data-Juicer 2.0 has been encapsulated into numerous effect-proven and illustrative data recipes (in YAML format as shown in Fig. 4), catering to various vertical domains, such as multimodal contrastive data synthesis (Jiao et al., 2024) and persona-oriented data processing, maintained within the DJ cookbook online. These are user-friendly, readily deployable, and facilitate interface exposure and reuse across different levels (Sec. 4), simplifying recipe routing and tailored processing solution generation by foundation-model-based agents.
@@
To avoid the substantial costs from trial and error for data and model development in foundation model scenarios, we further develop a Sandbox suite in Data-Juicer 2.0 for data-model co-development, serving as a specialized intermediate layer connecting data processing jobs to numerous open-source infrastructures of model training and evaluation. The suite offers template workflows that extend beyond dataset-only development by incorporating cost-effective model training and evaluation signals, and quantitatively studying interplays among data, model, and compute. Users can easily conduct small, quick, and comparative experiments to find insights and superior data recipes, which can then be scaled to larger models and datasets, thereby optimizing the return on investment in data-model co-development. Additional details and empirically validated support can be found in (Chen et al., 2024c).
5.4. Internal Adaptation for DJ-Dataset
5.4.1. Engine-Agnostic Processing

### v2

@@ D.3 Web & CMD tools
Thanks to the low-level and RESTful APIs, we further establish service-level capabilities across various scenarios in Data-Juicer 2.0, encapsulating scenario-specific processing solutions into built-in data recipes, described as YAML configuration files for end-to-end workflows (an example for video-data-synthesis is shown in Fig. 6). Centered around these data recipes, we provide user-friendly, highly encapsulated command-line interfaces as illustrated in Listing D.3, supporting automated lightweight installation and recipe-level invocation. These tools automatically gather and install the requisite of OPs within the recipe, perform automated OP fusion, and maximize adaptive utilization of computing resources while minimizing user cognitive load and operational cost.
@@
Utilizing the aforementioned APIs and tools, Data-Juicer’s capabilities are seamlessly integrated into Alibaba Cloud’s visual modeling product, PAI-Designer [18]. As shown in Fig. 1, Data-Juicer-Operators are encapsulated as visual components within Designer, presented in a directory tree structure for easy access. Users can manage data by intuitive drag-and-drop manners, and import built-in or customized recipes in YAML format using graphical canvas interface. With Designer as entry point, users can activate MaxCompute cluster nodes [17] or containerized deep compute clusters (PAI-DLC) [16] directly for efficient cloud-native distributed data processing, leveraging Alibaba Cloud’s data storage solutions (e.g., Object Storage OSS [20], Cloud Parallel File System CPFS [15], and NAS [19]) along with automated performance monitoring and optimization.
D.4 Natural Language Interaction
User: Filter out too short text samples. The dataset path is input_data.jsonl Assistant: <thought>Filter out text samples that are too short based on the provided dataset path.</thought>
@@ D.4 Natural Language Interaction
Assistant: <thought>The filtering has been successfully executed, and the filtered dataset is now at the specified path.</thought>
@@
Assistant: Your request to filter out too short text samples has been completed. The processed dataset is available at output.jsonl.
Querying and understanding over 100 diverse OPs have become challenging. Thankfully, advances in foundation models have transformed human-computer interactions. Research shows the synergy of reasoning and acting (ReAct) [77] in language models enhances task-solving. Using RESTful API services, we developed model-driven agents based on ReAct, enabling intuitive data processing through natural language. We adopted AgentScope [28], a multi-agent platform, for low-code integration, using prompt-optimized ReAct agents to align OP functionalities with our RESTful APIs. Built-in function preprocessing, response parsing, and fault-tolerance streamline development. Data-Juicer 2.0 thus allows flexible instruction input, interpreting user intents to execute appropriate data processing functions.
@@
Appendix E Core Runtime Implementation Details
E.1 Execution Mode Configuration
The core runtime of Data-Juicer 2.0 is implemented in Python, featuring a flexible data processing pipeline characterized by two execution modes: a standalone mode for convenient single-node execution, based on the Hugging Face Dataset [36], and a distributed mode that offers scalability across multiple nodes, leveraging the Ray Data [45] and MaxFrame-DataFrame [17]. These frameworks offer diverse computational engines, each with distinct capabilities and suitable use cases, but they also come with heterogeneous programming interfaces and complex implementation details.
In Data-Juicer 2.0, we exploit the strengths of these diverse computational engines and their associated programming classes by abstracting a top-level Data-Juicer-Dataset class and devising a standardized data representation schema (detailed in the next subsection). The primary design principle employed is the Facade Pattern. The abstracted class provides comprehensive and unified development interfaces, facilitating seamless use across standalone, distributed, and cloud cluster environments, meanwhile optimized to support extensive multimodal data processing.
@@ E.1 Execution Mode Configuration
data.export(tgt_file) # tgt_file can be either local path or remote path Some basic interfaces of Data-Juicer-Dataset.
Moreover, following the principle of Template Method Pattern, the Data-Juicer-Dataset class establishes templated workflows for OP executing on datasets and provides several basic and unified interfaces, catering to users with diverse programming preferences, as demonstrated in Listing E.1. The complexities of underlying computational engines and cumbersome runtime issues related to instantiated dataset objects are abstracted away from these interfaces, ensuring transparency and ease of use for end-users and developers alike.
New functionalities can be added by merely implementing new OP classes without altering the internal logic of Data-Juicer-Dataset. The standalone and distributed execution modes automatically recognize and switch configurations based on runtime settings. Coupled with unified function signatures, the data processing can be consolidated into a batched processing mode, not only standardizing the data interface for OP processing but also facilitating the implementation of a robust sample-level fault tolerance mechanism. The internal adaptation features are elaborated in Appendix F.2.
E.2 Data Schema Implementation
@@
[!h] {minted} [ frame=lines, framesep=2mm, baselinestretch=1.2, fontsize=, linenos, breaklines=true, bgcolor=lightgray!10, rulecolor=black, numbersep=5pt, ] json // >>> core contents: texts, dialogs, … // for general pretraining "text": "<__dj__image> desc of image 1 <|__dj__eoc|> desc of image 2 <__dj__image> <|__dj__eoc|>", // for post-tuning "query": "<query2>", "response": "<response2>", "history": [["<query1>", "<response1>"]], // <<< core contents
// $>$ extra data contents: multimodal data paths, … "images": [ "path/to/the/image1", "path/to/the/image2" ], // $<$ extra data contents
// $>$ meta infos and stats, which could be primitive or produced by Data-Juicer "meta": "src": "customized", "version": "0.1" , "stats": "lang": "en", "image_widths": [224, 336] , // $<$ meta infos and stats Illustration of data schema of Data-Juicer 2.0.
@@
In the context of foundation models and post-tuning tasks, datasets are complex, originating from various sources and modalities. Target tasks are diverse, requiring different data formats and organizational structures. Integrating diverse data processing into a unified pipeline is challenging. Data-Juicer 2.0 adopts an extensible and intermediate data representation schema that intrinsically supports multimodal data types and flexible data field customization.
@@
An example of a data sample is illustrated in Listing E.2. In the data schema, each dataset encapsulates column-stored samples represented by a non-recursive dictionary whose level-1 fields include three categories: (1) the core contents within customizable fields, including “text” that is usually used in the pretraining task, and “query”, “response”, “history” fields represent dialogs in post-tuning tasks, which are directly related to the pretraining or post-tuning procedures in the downstream usage of the dataset; (2) the extra data contents contain the path lists of multimodal data for multimodal datasets; (3) the “meta” information concerning this sample that stems from either its raw data or Data-Juicer’s Mappers for tagging, and the “stats” information calculated by Data-Juicer’s Filters (all Filter OPs consist of a compute_stats() method, utilized for subsequent analysis and conditional filtering).
@@
For multimodal datasets, the content is primarily centered on the textual modality within the “text” field, capable of optionally including several text chunks split by customizable end-of-chunk tokens, which are <|__dj__eoc|> in default. Each chunk serves as a semantic unit, where the associated multimodal data within the same chunk relates to the same topic, thereby aligning with one another. Multimodal data (excluding text) are denoted by ordered customized special tokens in the text (e.g., <__dj__image>), storing entities at designated files, accessible through indexed file paths within the modality fields of samples to facilitate object sharing and elimination of redundant computing. For instance, in the example shown in Listing E.2, the first chunk in the “text” field split by the <__dj__eoc> token contains the first <__dj__image> token and a description to this image, which corresponds to the first image stored in the “images” field.
@@
This token-centric representation elegantly aligns multimodal data while preserving positional information, friendly to current prevalent learning paradigms for foundation models employing next-token prediction tasks [5, 72]. And it’s compatible with both simple cross-modal pairing datasets (e.g., image-text pairs [14] and video-text pairs [75]), as well as complex interleaved multimodal datasets (e.g., MMC4 [83]), owing to its chunkable and special token design.
@@
For post-tuning datasets, we provide several core fields in the core contents to represent the dialog datasets in our intermediate format, which is naturally aligned with Alpaca and Query-Response formats in the ModelScope-SWIFT repository [80] for smooth training of 400+ foundation models. As for other widely-used formats in well-known repositories like LLaMA-Factory [81], a suite of dataset conversion tools was developed to efficiently transform various popular multimodal datasets in them to and from the Data-Juicer schema, enabling users to process extensive existing models and datasets within Data-Juicer. These tools also serve as demonstrations for extending support to other uncovered datasets.
E.3 Operator Factory & Taxonomy
E.3.1 OP Design Principles
@@
Operators are fundamental units responsible for executing processing functionalities, such as enhancing the accuracy of descriptive text or filtering out images containing Not-Safe-for-Work (NSFW) content. As illustrated by the middle orange box in Fig. 1, Data-Juicer defines five previous atomic OP classes: Formatter, Filter, Mapper, Deduplicator, and Selector; alongside four new types of compositional OPs: Grouper, Aggregator, FusedOP, and ScriptOP. The first five OP classes handle dataset format conversion, sample filtering, modification, deduplication, and selection, respectively. Following the Strategy Pattern, they are designed to encapsulate diverse algorithms that can be dynamically used to process the data. Each OP has a clearly defined role and can be interchanged or modified without impacting other system parts.
@@
Moreover, following the Decorator Pattern, compositional OPs are provided to enhance existing functionality without modifying the prevailing OP structure while dynamically adding data processing behavior to objects. The FusedOP enables explicitly grouping multiple atomic OPs to process data in a fine-grained manner within the same data batch, opposed to the default sequential processing across datasets as demonstrated by Listing E.3.1. The Grouper takes a Data-Juicer-Dataset as input and groups data samples into batches, which can then be input into the Aggregator for subsequent aggregation. For example, we can employ extract_entity_attribute_mapper and entity_attribute_aggregator following a key_value_grouper for meta-information extraction from textual input. Meanwhile, the ScriptOP allows users to incorporate existing Python files or execute code snippets, utilizing customized functionalities encoded within scripts like helper_func.py. This includes leveraging existing Data-Juicer command tools or integrating short Python scripts (e.g., lambda functions).
@@
E.3.2 Unified Coordination of Logical Operations
@@
Specifically, a top-level OP factory class is abstracted above the aforementioned fundamental OP classes. In this class, functionalities common to all OPs are extracted, such as preprocessing of instantiated parameters, support for serialization, and configuration of OP-aware runtime parallelism. Besides, a unified run() method is implemented, maintaining a consistent interface for integration and API calls. Furthermore, parallelism in multi-processing and multi-GPU is automatically configured and decoupled from specific OPs, ensuring transparency for end users and developers, as introduced subsequently in Appendix G. Lower-level OP classes define their own templated execution logic behind the run invocation. Taking the Filter class as an example, its core logic first engages the compute_stats() method to obtain statistical information based on specific metrics and then invokes its process() method to determine sample filtering based on thresholds.
@@ E.4 Control Panel Implementation
With the fundamental Data-Juicer-Dataset and Data-Juicer-Operators primitives established, we offer a series of control panel modules (the red box in Fig. 1) to organically combine them and accomplish data processing tasks. Executor encapsulates a series of standardized execution processes tailored for different standalone and distributed engines, leveraging modules such as Config and Monitor to accomplish end-to-end system configuration, data loading, analysis, processing iteration, data checkpointing, and more. Template workflows manage the complete data development lifecycle, including feedback, data storage, logging, and performance and operational monitoring.
@@
To avoid the substantial costs from trial and error for data and model development in foundation model scenarios, we further develop a Sandbox suite in Data-Juicer 2.0 for data-model co-development, serving as a specialized intermediate layer connecting data processing jobs to numerous open-source infrastructures of model training and evaluation. The suite offers template workflows that extend beyond dataset-only development by incorporating cost-effective model training and evaluation signals, and quantitatively studying interplays among data, model, and compute. Users can easily conduct small, quick, and comparative experiments to find insights and superior data recipes, which can then be scaled to larger models and datasets, thereby optimizing the return on investment in data-model co-development. Additional details and empirically validated support can be found in [13].
Appendix F Optimization Details on Data-Juicer-Dataset
F.1 Engine-agnostic Processing